# Permutation polynomials with a few terms over finite fields

### v2

@@ Abstract.
The work of S. Hasan was supported by the Science and Engineering Research Board, Government of India under CRG/2022/005418. The work of C. Li was supported by the Research Council of Norway under Grant No. 311646. The work of M. Pal was supported by the Research Council of Norway under Grant No. 314395.
@@ 1. Introduction
Permutation polynomials with a few terms are of particular interest due to their simple algebraic structures, which make them suitable for implementation in various applications. A lot of research has been done in recent years on the construction of permutation polynomials with a few terms. The most simple polynomials are the monomials $X^{n}$, which permutes $\mathbb{F}_{q}$ if and only if $\gcd(n,q-1)=1$. This fact makes the classification of permutation monomials easy. On the other hand, the classification of permutation polynomials having a few terms, such as binomials, trinomials, quadrinomials, and pentanomials, is non-trivial and has not yet been completely resolved. In order to study permutation binomials, Hou argued in [22] that it is enough to consider binomials of the form $f(X)=X^{r}(X^{\frac{q-1}{d}}+a)$, where $r\geq 1,d\mid(q-1)$ and $a\in\mathbb{F}_{q}^{*}$. For $r=1$ and a fixed $d$, Carlitz and Wells  [8] proved that there always exist elements $a\in\mathbb{F}_{q}^{*}$ such that $f(X)$ permutes $\mathbb{F}_{q}$ for sufficiently large $q$. A lot of research [27, 38, 41, 42] has been done on the characterization of $r’$ ’s and $a’$ ’s, for which $f(X)$ is a permutation binomial. For binomials of the form $f(X)=X^{n}+aX^{m}$ Turnwald in [52], and further, Masuda and Zieve in [41] have given certain conditions on the positive integers $m,n$, and $a\in\mathbb{F}_{q}^{*}$ for which $f(X)$ are not permutations of $\mathbb{F}_{q}$. Further, the permutation binomials having the form $X(X^{r(q-1)}+a)$ over $\mathbb{F}_{q^{2}}$ for $r\in\{2,3,5,7\}$ are investigated in  [20, 25, 30]. For explicit constructions of permutation binomials and trinomials so far, the interested reader may refer to  [4, 6, 19, 26, 31, 33, 53].
@@ 2. Preliminaries
A function of the form $f(X)=\frac{P(X)}{Q(X)}$ defines a mapping from $\mathbb{P}^{1}(\mathbb{F}_{q})=\mathbb{F}_{q}\cup\{\infty\}$ to itself, where $P(X),Q(X)\in\mathbb{F}_{q}[X]$ and $\gcd(P,Q)=1$, is called rational function. We denote the degree of $f$ by $\deg(f)$, i.e.,$\deg(f):=\max(\deg(P),\deg(Q))$. The set of all such rational functions is denote by $\mathbb{F}_{q}(X)$. We call $f(X)$ a permutation rational (PR) function if the induced map $X\mapsto f(X)$ is a bijection from $\mathbb{P}^{1}(\mathbb{F}_{q})$ to itself.
@@
It is easy to see that for equivalent functions $f,g\in\mathbb{F}_{q}(X)$, $f$ permutes $\mathbb{P}^{1}(\mathbb{F}_{q})$ if and only if $g$ permutes $\mathbb{P}^{1}(\mathbb{F}_{q})$. We now recall some lemmas that give us the complete classification of permutation rational functions of degree at most four over $\mathbb{F}_{q}$.
@@
    $q\leq 8$ and $f(X)$ is equivalent to a rational function in Table $1$ [13, Page 19]
@@
Two permutation polynomials $f(X)$ and $g(X)$ in $\mathbb{F}_{q}[X]$ are called quasi-multiplicative equivalent if there exists an integer $11\leq d<q-$ 1 with $\gcd(d,q-1)=1$ and $f(X)=ug(vX^{d})$, where $u,v\in\mathbb{F}_{q}^{*}$.
@@ 3. Permutation polynomials from degree-three rational functions
$\rho(X):=\frac{\delta X-\beta\delta^{q}}{X-\beta},~{}\mbox{and}~{}{\nu}(X):={\tilde{\beta}}\left(\frac{X-{\tilde{\delta}}^{q}}{X-{\tilde{\delta}}}\right),$
@@
$\begin{split}({\nu}\circ X^{3}\circ\rho)(X)&={\nu}\circ\left(\frac{\delta X-\beta\delta^{q}}{X-\beta}\right)^{3}\\ &={\tilde{\beta}}\left(\frac{(\delta X-\beta\delta^{q})^{3}-{\tilde{\delta}}^{q}(X-\beta)^{3}}{(\delta X-\beta\delta^{q})^{3}-{\tilde{\delta}}(X-\beta)^{3}}\right)\\ &={\tilde{\beta}}\left(\frac{(\delta^{3}-{\tilde{\delta}}^{q})X^{3}-3\beta(\delta^{q+2}-{\tilde{\delta}}^{q})X^{2}+3\beta^{2}(\delta^{2q+1}-{\tilde{\delta}}^{q})X-\beta^{3}(\delta^{3q}-{\tilde{\delta}}^{q})}{(\delta^{3}-{\tilde{\delta}})X^{3}-3\beta(\delta^{q+2}-{\tilde{\delta}})X^{2}+3\beta^{2}(\delta^{2q+1}-{\tilde{\delta}})X-\beta^{3}(\delta^{3q}-{\tilde{\delta}})}\right)\\ &=\frac{N_{3}X^{3}+N_{2}X^{2}+N_{1}X+N_{0}}{D_{3}X^{3}+D_{2}X^{2}+D_{1}X+D_{0}},\end{split}$
@@
(3.1) $\begin{cases}N_{0}&=-{\tilde{\beta}}\beta^{3}(\delta^{3q}-{\tilde{\delta}}^{q}),\\ N_{1}&=3{\tilde{\beta}}\beta^{2}(\delta^{2q+1}-{\tilde{\delta}}^{q}),\\ N_{2}&=-3{\tilde{\beta}}\beta(\delta^{q+2}-{\tilde{\delta}}^{q}),\\ N_{3}&={\tilde{\beta}}(\delta^{3}-{\tilde{\delta}}^{q}),\end{cases}~{}~{}~{}\mbox{and}~{}~{}~{}\begin{cases}D_{0}&=-\beta^{3}(\delta^{3q}-{\tilde{\delta}}),\\ D_{1}&=3\beta^{2}(\delta^{2q+1}-{\tilde{\delta}}),\\ D_{2}&=-3\beta(\delta^{q+2}-{\tilde{\delta}}),\\ D_{3}&=(\delta^{3}-{\tilde{\delta}}).\end{cases}$
@@
Let $q\equiv 2\pmod{3}$ and $h(X)=D_{3}X^{3}+D_{2}X^{2}+D_{1}X+D_{0}$, where $D_{0},D_{1},D_{2},$ $D_{3}$ are as given in Equation (3.1), and $\beta,{\tilde{\beta}}$ be elements in $\mu_{q+1}$ satisfying $01+{\tilde{\beta}}\beta^{3}=$ 0. Moreover, $\delta$ and $\tilde{\delta}$ in the expressions of $D_{i}$’s, $0\leq i\leq 3$ are such that $\delta,\tilde{\delta}\in\mathbb{F}_{q^{2}}\setminus\mathbb{F}_{q}$ and $\tilde{\delta}\not\in\{\delta^{3},\delta^{q+2},\delta^{2q+1},\delta^{3q}\}$. Then
@@
From Lemma 2.7, we know that $f$ is a permutation of $\mathbb{F}_{q^{2}}$ if and only if $\gcd(3,q-1)=1$ and $g(X):=X^{3}h(X)^{q-1}$ permutes $\mu_{q+1}$. Since $q\equiv 2\pmod{3}$, the first condition holds trivially. Therefore it is sufficient to show that if $01+{\tilde{\beta}}\beta^{3}=$ 0 then $g$ permutes $\mu_{q+1}$. From Remark 3.1, we know that $h(X)=0$ has no solution in $\mu_{q+1}$. Now for any $\alpha\in\mu_{q+1}$, consider
@@
where $N_{0},N_{1},N_{2}$ and $N_{3}$ are as given in Equation (3.1). If all the four conditions in the system (3.2) are satisfied then ${\tilde{g}}={\nu}\circ X^{3}\circ\rho$. It is easy to verify that if $01+{\tilde{\beta}}\beta^{3}=$ 0 then all the four conditions of the system (3.2) are satisfied. Thus, if $01+{\tilde{\beta}}\beta^{3}=$ 0 then ${\tilde{g}}$ is a permutation of $\mu_{q+1}$ and consequently $f$ is a permutation of $\mathbb{F}_{q^{2}}$. ∎
@@
Let $q\equiv 2\pmod{3}$ and $D_{0},D_{1},D_{2},D_{3}$ are as given in the Equation (3.1). Moreover, $\delta$ and $\tilde{\delta}$ in the expressions of $D_{i}$’s, $0\leq i\leq 3$ are such that $\delta,\tilde{\delta}\in\mathbb{F}_{q^{2}}\setminus\mathbb{F}_{q}$ and $\tilde{\delta}\not\in\{\delta^{3},\delta^{q+2},\delta^{2q+1},\delta^{3q}\}$. If $01+{\tilde{\beta}}\beta^{3}=$ 0 for some ${\tilde{\beta}}\in\mu_{q+1}$ then the following quadrinomials are permutation quadrinomials over $\mathbb{F}_{q^{2}}$
@@
$\rho(X)=\frac{\delta X-\beta\delta^{q}}{X-\beta},~{}\mbox{and}~{}{\nu}(X)={\tilde{\beta}}\left(\frac{X-{\tilde{\delta}}^{q}}{X-{\tilde{\delta}}}\right),$
@@
(3.3) $\begin{cases}N_{0}&={\tilde{\beta}}\beta^{3}(\alpha\delta^{q}-\delta^{3q}+{\tilde{\delta}}^{q}),\\ N_{1}&=\alpha{\tilde{\beta}}\beta^{2}(\delta^{q}-\delta),\\ N_{2}&=\alpha{\tilde{\beta}}\beta(\delta^{q}-\delta),\\ N_{3}&={\tilde{\beta}}(\delta^{3}-\alpha\delta-{\tilde{\delta}}^{q}),\end{cases}~{}~{}~{}\mbox{and}~{}~{}~{}\begin{cases}D_{0}&=\beta^{3}(\alpha\delta^{q}-\delta^{3q}+{\tilde{\delta}}),\\ D_{1}&=\alpha\beta^{2}(\delta^{q}-\delta),\\ D_{2}&=\alpha\beta(\delta^{q}-\delta),\\ D_{3}&=\delta^{3}-\alpha\delta-{\tilde{\delta}},\end{cases}$
@@
Let $q\equiv 0\pmod{3}$ and $h(X)=D_{3}X^{3}+D_{2}X^{2}+D_{1}X+D_{0}$, where $D_{0},D_{1},D_{2},$ $D_{3}$ are given in Equation (3.3). Moreover, $\delta$ and $\tilde{\delta}$ in the expressions of $D_{i}$’s, $0\leq i\leq 3$ are such that $\delta,\tilde{\delta}\in\mathbb{F}_{q^{2}}\setminus\mathbb{F}_{q}$ and $\tilde{\delta}\not\in\{\delta^{3q}-\alpha\delta^{q},\delta^{3}-\alpha\delta\}$, where $\alpha$ is either zero or a non-square in $\mathbb{F}_{q}$. Let $\beta,{\tilde{\beta}}$ be elements in $\mu_{q+1}$ satisfying $01+{\tilde{\beta}}\beta^{3}=$ 0. Then $f(X)=X^{3}h(X^{q-1})$ is a PP over $\mathbb{F}_{q^{2}}$.
@@
Let $q\equiv 0\pmod{3}$ and $D_{0},D_{1},D_{2},$ $D_{3}$ are given in Equation (3.3). Moreover, $\delta$ and $\tilde{\delta}$ in the expressions of $D_{i}$’s, $0\leq i\leq 3$ are such that $\delta,\tilde{\delta}\in\mathbb{F}_{q^{2}}\setminus\mathbb{F}_{q}$ and $\tilde{\delta}\not\in\{\delta^{3q}-\alpha\delta^{q},\delta^{3}-\alpha\delta\}$, where $\alpha$ is either zero or a non-square in $\mathbb{F}_{q}$. If $01+{\tilde{\beta}}\beta^{3}=$ 0 for some ${\tilde{\beta}}\in\mu_{q+1}$ then the following quadrinomials are permutation quadrinomials over $\mathbb{F}_{q^{2}}$

### v3

@@ Abstract.
The work of S. U. Hasan was supported by the Science and Engineering Research Board (SERB), Government of India (Grant No. CRG/2022/005418), and the Indo-Norwegian Cooperation Programme 2024 (Project No. INCP2-2024/10213). He also acknowledges the support of SPIRE project from Univ. of Bergen, Norway. H. Kumar acknowledges support from the Prime Minister’s Research Fellowship (PMRF), Government of India, under PMRF ID 3002900 at IIT Jammu. C. Li’s work was supported by the Research Council of Norway (Grant No. 311646) and the Indo-Norwegian Cooperation Programme 2024 (Project No. INCP2-2024/10213). M. Pal acknowledges support from the Research Council of Norway under Grant No. 314395.
@@ 1. Introduction
Permutation polynomials with a few terms are of particular interest due to their simple algebraic structures, which make them suitable for implementation in various applications. A lot of research has been done in recent years on the construction of permutation polynomials with a few terms. The most simple polynomials are the monomials $X^{n}$, which permutes $\mathbb{F}_{q}$ if and only if $\gcd(n,q-1)=1$. This fact makes the classification of permutation monomials easy. On the other hand, the classification of permutation polynomials having a few terms, such as binomials, trinomials, quadrinomials, and pentanomials, is non-trivial and has not yet been completely resolved. In order to study permutation binomials, Hou argued in [22] that it is enough to consider binomials of the form $f(X)=X^{r}(X^{\frac{q-1}{d}}+a)$, where $r\geq 1,d\mid(q-1)$ and $a\in\mathbb{F}_{q}^{*}$. For $r=1$ and a fixed $d$, Carlitz and Wells  [8] proved that there always exist elements $a\in\mathbb{F}_{q}^{*}$ such that $f(X)$ permutes $\mathbb{F}_{q}$ for sufficiently large $q$. A lot of research [27, 38, 41, 42] has been done on the characterization of $r$’s and $a$’s, for which $f(X)$ is a permutation binomial. For binomials of the form $f(X)=X^{n}+aX^{m}$ Turnwald in [52], and further, Masuda and Zieve in [41] have given certain conditions on the positive integers $m,n$, and $a\in\mathbb{F}_{q}^{*}$ for which $f(X)$ are not permutations of $\mathbb{F}_{q}$. Further, the permutation binomials having the form $X(X^{r(q-1)}+a)$ over $\mathbb{F}_{q^{2}}$ for $r\in\{2,3,5,7\}$ are investigated in  [20, 25, 30]. For explicit constructions of permutation binomials and trinomials so far, the interested reader may refer to  [4, 6, 19, 26, 31, 33, 53].
@@ 2. Preliminaries
[…]
@@
It is easy to see that for equivalent functions $g\in\mathbb{F}_{q}(X), ff$ permutes $\mathbb{P}^{1}(\mathbb{F}_{q})$ if and only if $g$ permutes $\mathbb{P}^{1}(\mathbb{F}_{q})$. We now recall some lemmas that give us the complete classification of permutation rational functions of degree at most four over $\mathbb{F}_{q}$.
@@
    $q\leq 8$ and $f(X)$ is equivalent to a rational function in Table 11 [13, Page 19]
@@
Two permutation polynomials $f(X)$ and $g(X)$ in $\mathbb{F}_{q}[X]$ are called quasi-multiplicative equivalent if there exists an integer $1\leq d<q-1$ with $\gcd(d,q-1)=1$ and $f(X)=ug(vX^{d})$, where $u,v\in\mathbb{F}_{q}^{*}$.
@@ 3. Permutation polynomials from degree-three rational functions
[…]
@@
$\begin{split}&{\nu}\circ\left(\frac{\delta X-\beta\delta^{q}}{X-\beta}\right)^{3}\\ &={\tilde{\beta}}\left(\frac{(\delta X-\beta\delta^{q})^{3}-{\tilde{\delta}}^{q}(X-\beta)^{3}}{(\delta X-\beta\delta^{q})^{3}-{\tilde{\delta}}(X-\beta)^{3}}\right)\\ &={\tilde{\beta}}\left(\frac{(\delta^{3}-{\tilde{\delta}}^{q})X^{3}-3\beta(\delta^{q+2}-{\tilde{\delta}}^{q})X^{2}+3\beta^{2}(\delta^{2q+1}-{\tilde{\delta}}^{q})X-\beta^{3}(\delta^{3q}-{\tilde{\delta}}^{q})}{(\delta^{3}-{\tilde{\delta}})X^{3}-3\beta(\delta^{q+2}-{\tilde{\delta}})X^{2}+3\beta^{2}(\delta^{2q+1}-{\tilde{\delta}})X-\beta^{3}(\delta^{3q}-{\tilde{\delta}})}\right)\\ &=\frac{N_{3}X^{3}+N_{2}X^{2}+N_{1}X+N_{0}}{D_{3}X^{3}+D_{2}X^{2}+D_{1}X+D_{0}},\end{split}$
@@
[…]
@@
Let $q\equiv 2\pmod{3}$ and $h(X)=D_{3}X^{3}+D_{2}X^{2}+D_{1}X+D_{0}$, where $D_{0},D_{1},D_{2},$ $D_{3}$ are as given in Equation (3.1), and $\beta,{\tilde{\beta}}$ be elements in $\mu_{q+1}$ satisfying $1+{\tilde{\beta}}\beta^{3}=0$. Moreover, $\delta$ and $\tilde{\delta}$ in the expressions of $D_{i}$’s, $0\leq i\leq 3$ are such that $\delta,\tilde{\delta}\in\mathbb{F}_{q^{2}}\setminus\mathbb{F}_{q}$ and $\tilde{\delta}\not\in\{\delta^{3},\delta^{q+2},\delta^{2q+1},\delta^{3q}\}$. Then
@@
From Lemma 2.7, we know that $f$ is a permutation of $\mathbb{F}_{q^{2}}$ if and only if $\gcd(3,q-1)=1$ and $g(X):=X^{3}h(X)^{q-1}$ permutes $\mu_{q+1}$. Since $q\equiv 2\pmod{3}$, the first condition holds trivially. Therefore it is sufficient to show that if $1+{\tilde{\beta}}\beta^{3}=0$ then $g$ permutes $\mu_{q+1}$. From Remark 3.1, we know that $h(X)=0$ has no solution in $\mu_{q+1}$. Now for any $\alpha\in\mu_{q+1}$, consider
@@
where $N_{0},N_{1},N_{2}$ and $N_{3}$ are as given in Equation (3.1). If all the four conditions in the system (3.2) are satisfied then ${\tilde{g}}={\nu}\circ X^{3}\circ\rho$. It is easy to verify that if $1+{\tilde{\beta}}\beta^{3}=0$ then all the four conditions of the system (3.2) are satisfied. Thus, if $1+{\tilde{\beta}}\beta^{3}=0$ then $\tilde{g}$} is a permutation of $\mu_{q+1}$ and consequently $f$ is a permutation of $\mathbb{F}_{q^{2}}$. ∎
Example 3.4.
Let $q=5$, $g$ be a generator of the cyclic group $\mathbb{F}_{q^{2}}^{*}$, $\beta=-1$, $\tilde{\beta}=1$ and $\delta=g=\tilde{\delta}$. From Theorem  3.3, we obtain a permutation quadrinomial $f(X)=3(g+1)X^{15}+3gX^{11}+(g+1)X^{7}+2X^{3}$ of $\mathbb{F}_{q^{2}}$. In addition, let $q=5^{3}$, $g$ be a primitive element of $\mathbb{F}_{q^{2}}$ satisfying $g^{6}+g^{4}+g^{3}+g^{2}+2=0$, $(\beta,\tilde{\beta})=(-1,1)$ and $(\delta,\tilde{\delta})=(g^{14078},g^{6470})$. Then from Theorem  3.3, we get a permutation quadrinomial $f(X)=g^{12017}X^{3q}+g^{9477}X^{2q+1}+g^{10055}X^{q+2}+g^{7976}X^{3}$ of $\mathbb{F}_{q^{2}}$.
@@
Let $q\equiv 2\pmod{3}$ and $D_{0},D_{1},D_{2},D_{3}$ are as given in the Equation (3.1). Moreover, $\delta$ and $\tilde{\delta}$ in the expressions of $D_{i}$’s, $0\leq i\leq 3$ are such that $\delta,\tilde{\delta}\in\mathbb{F}_{q^{2}}\setminus\mathbb{F}_{q}$ and $\tilde{\delta}\not\in\{\delta^{3},\delta^{q+2},\delta^{2q+1},\delta^{3q}\}$. If $1+{\tilde{\beta}}\beta^{3}=0$ for some ${\tilde{\beta}}\in\mu_{q+1}$ then the following quadrinomials are permutation quadrinomials over $\mathbb{F}_{q^{2}}$
@@
Example 3.6.
Consider $q=2^{3}$ and $\mathbb{F}_{q^{2}}^{*}=\langle g\rangle$, where $g^{6}+g^{4}+g^{3}+g+1=0$. Let $\beta=1=\tilde{\beta}$ and $\delta=g=\tilde{\delta}$. Then Theorem  3.5 (1) implies that $f(X)=(g^{5}+g^{3}+1)X^{57}+(g^{3}+g)X^{15}+(g^{5}+g^{4}+g+1)X^{8}+(g^{5}+g^{2})X$ is a permutation quadrinomial over $\mathbb{F}_{q^{2}}$.
@@
[…]
@@
[…]
@@
Let $q\equiv 0\pmod{3}$ and $h(X)=D_{3}X^{3}+D_{2}X^{2}+D_{1}X+D_{0}$, where $D_{0},D_{1},D_{2},$ $D_{3}$ are given in Equation (3.3). Moreover, $\delta$ and $\tilde{\delta}$ in the expressions of $D_{i}$’s, $0\leq i\leq 3$ are such that $\delta,\tilde{\delta}\in\mathbb{F}_{q^{2}}\setminus\mathbb{F}_{q}$ and $\tilde{\delta}\not\in\{\delta^{3q}-\alpha\delta^{q},\delta^{3}-\alpha\delta\}$, where $\alpha$ is either zero or a non-square in $\mathbb{F}_{q}$. Let $\beta,{\tilde{\beta}}$ be elements in $\mu_{q+1}$ satisfying $1+{\tilde{\beta}}\beta^{3}=0$. Then $f(X)=X^{3}h(X^{q-1})$ is a PP over $\mathbb{F}_{q^{2}}$.
Example 3.10.
In the above Theorem 3.9, if we choose $\beta=2$, $\tilde{\beta}=1$, $\delta=g=\tilde{\delta}$ and $\alpha=2$. Then we get $f(X)=(g+1)X^{7}+2(g+1)X^{5}+(2g+1)X^{3}+(2g+1)X$, a permutation polynomial over the finite field $\mathbb{F}_{q^{2}}$, where $q=3$ and $g$ is generator of the cyclic group $\mathbb{F}_{q^{2}}^{*}$. In addition, by taking $q=3^{4}$, a primitive element $g$ of $\mathbb{F}_{q^{2}}$ satisfying $g^{8}+2g^{5}+g^{4}+2g^{2}+2g+2=0$, $\alpha=g^{q+1}$, $(\beta,\beta_{1})=(-1,1)$ and $(\delta,\tilde{\delta})=(g^{4898},g^{332})$, we can obtain a permutation quadrinomial $f(X)=g^{1523}X^{3q}+g^{1681}X^{2q+1}+g^{4961}X^{q+2}+g^{3736}X^{3}$ over $\mathbb{F}_{q^{2}}$.
@@
Let $q\equiv 0\pmod{3}$ and $D_{0},D_{1},D_{2},$ $D_{3}$ are given in Equation (3.3). Moreover, $\delta$ and $\tilde{\delta}$ in the expressions of $D_{i}$’s, $0\leq i\leq 3$ are such that $\delta,\tilde{\delta}\in\mathbb{F}_{q^{2}}\setminus\mathbb{F}_{q}$ and $\tilde{\delta}\not\in\{\delta^{3q}-\alpha\delta^{q},\delta^{3}-\alpha\delta\}$, where $\alpha$ is either zero or a non-square in $\mathbb{F}_{q}$. If $1+{\tilde{\beta}}\beta^{3}=0$ for some ${\tilde{\beta}}\in\mu_{q+1}$ then the following quadrinomials are permutation quadrinomials over $\mathbb{F}_{q^{2}}$
@@
Example 3.12.
Let $q=3^{2}$ and $g$ be a generator of the cyclic group $\mathbb{F}_{q^{2}}^{*}$ satisfying $X^{4}+2X^{3}+2=0$. Consider $\beta=1$, $\tilde{\beta}=-1$, $\delta=g=\tilde{\delta}$ and $\alpha=g^{q+1}$. Then we obtain $f(X)=(g+1)X^{27}+(g^{3}+g+1)X^{19}+(g^{3}+g+1)X^{11}+(g^{3}+g^{2}+2)X^{3}$, a permutation quadrinomial over the finite field $\mathbb{F}_{q^{2}}$ from Theorem 3.11 (3).